# Retinexmamba: Retinex-based Mamba for Low-light Image Enhancement

### v1

@@ 4.2 Low-light Image Enhancement
The qualitative comparison results between RetinexMamba and other SOTA algorithms are shown in Fig.4 and 5. Please zoom in for better visual clarity. Figure Three compares the LOLv1 dataset, while Fig. 5 compares the LOLv2_real dataset. As depicted in Figure Three, previous methods exhibit amplification of noise, such as in Retinex-Net, as well as cases of underexposure like KinD and overexposure like Uretinex-Net. Similarly, on the synthetic dataset in Fig. 5, KinD shows instances of overexposure in the air conditioner area at the top and underexposure in the staircase area at the bottom, while ZeroRCE exhibits extensive noise and artifacts. Likewise, in Retinexformer, underexposure around the bowling ball at the top of Fig. 4 and color distortion in the stadium at the bottom are observed. In contrast, our RetinexMamba effectively controls exposure intensity, reduces color distortion, and minimizes noise.

### v2

@@ 4.2 Low-light Image Enhancement
The qualitative comparison results between RetinexMamba and other SOTA algorithms are shown in Fig. 5 and 6. Please zoom in for better visual clarity. Figure Three compares the LOLv1 dataset, while Fig. 6 compares the LOLv2_real dataset. As depicted in Figure Three, previous methods exhibit amplification of noise, such as in Retinex-Net, as well as cases of underexposure like KinD and overexposure like Uretinex-Net. Similarly, on the synthetic dataset in Fig. 6, KinD shows instances of overexposure in the air conditioner area at the top and underexposure in the staircase area at the bottom, while ZeroRCE exhibits extensive noise and artifacts. Likewise, in Retinexformer, underexposure around the bowling ball at the top of Fig. 5 and color distortion in the stadium at the bottom are observed. In contrast, our RetinexMamba effectively controls exposure intensity, reduces color distortion, and minimizes noise. Additionally, we compared two images on the LOLv2_syn dataset in both 2D and 3D HSV color space in Figs. 7 and 4. The results demonstrate that the images enhanced by RetinexMamba are closest to the Ground Truth.
Figure 7: 2D image display of the HSV color space.